40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1078     Entry time: Thu Oct 23 20:47:28 2008
Author: pete 
Type: Configuration 
Category: PSL 
Subject: FSS LO calibration for MEDM 
Today I took a quick series of measurements to calibrate the FSS LO power measurement in the MEDM. This was done by using the spec.an. to measure the 21.5 MHz peak in dBm at the LO input to the FSS box on the PSL table, and recording the MEDM value, for attenuations applied at the FSS REF box output ranging from -5 dBm to -30 dBm.

I measured the loss due to the BNC cable I used, which was (19.66-19.50) dBm. I accounted for this and plotted ln(MEDM) vs. dBm on the attached plot. A linear fit of this gives the CALC field of a calc record for the IOC db:
6.29*LOGE(A)+5.36

Since no one knew how to do this nonlinear conversion in EPICS I will describe how to do it in detail tomorrow. It is simple, although it requires power cycling the scipe3 bunch (typing "reboot" or "ctl-x" at the command prompt took it down, but it did not come back). I did power cycle those computers a few times today.
Attachment 1: fss_lo_calibration.png  13 kB  | Hide | Hide all
fss_lo_calibration.png
ELOG V3.1.3-